the data below shows a tabulation on the production of a hypothetical product

output (Q) total cost

0 25

1 32

2 38

3 42

4 48

5 58

6 67

7 78

8 98

use the data to answer the following questions:-

  1. determine the total fixed cost
  2. average variable cost schedule
  3. marginal cost schedule
  4. suppose this product is produced on a perfect market and the price of the commodity= 10. Determine the output Q that will maximize the profits. What is the maximum profits achieved by the firm? (show all the workings)
1
Expert's answer
2021-07-13T12:23:07-0400

1. Determination of total fixed cost

Q TC FC

0 25 25

1 32 25

2 38 25

3 42 25

4 48 25

5 58 25

6 67 25

7 78 25

8 98 25

NOTE: Fixed costs do not change with change in output

2. Average variable cost (AVC) schedule:

TVC = TC -TFC

AVC = "\frac{TVC}{Q}"

Q TC TFC TVC AVC

0 25 25 0 0

1 32 2 7 7

2 38 25 13 6.5

3 42 25 17 5.67

4 48 25 23 5.75

5 58 25 33 6.6

6 67 25 42 7

7 78 25 53 7.57

8 98 25 73 9.13


3. Marginal cost schedule

MC = TC of the proceeding level minus TC of the preceding level

Q TC MC

0 25 0

1 32 7

2 38 6

3 42 4

4 48 6

5 58 10

6 67 9

7 78 11

8 98 20


4. Determination of profit-maximizing output:

NOTE: Profits are maximized at the point where Marginal Cost (MC) equals Marginal Revenue (MR).

If price (P) IS 10, TR (P"\times"Q) and MR will be:

Q TC MC TR MR

0 25 0 0 0

1 32 7 10 10

2 38 6 20 10

3 42 4 30 10

4 48 6 40 10

5 58 10 50 10

6 67 9 60 10

7 78 11 70 10

8 98 20 80 10


Therefore, the profit-maximizing output is 5units;

Maximum (economic) profit = TR - TC = 50-58 = -8 (negative economic profit)

NOTE: The firm has an incentive to exit the market since it's unprofitable to continue operating.